# Coherent feedback control for cavity optomechanical systems with a frequency-dependent mirror

### v2

@@ I Introduction
While the standard sideband cooling scheme [1] provides a powerful tool to access the mechanical ground state in optomechanical systems with high sideband resolution [14, 15, 16], in practice setups often are in the unresolved-sideband regime, meaning that simple sideband cooling no longer works. Quantum feedback has emerged as a candidate for ground-state cooling and control of optomechanical systems that are in the unresolved-sideband regime [17]. Compared to active feedback [18, 19, 20, 21, 22, 23, 24, 25], that has to face excess noise in the out-of-loop optical field and decoherence due to quantum measurement [22], coherent feedback is promising since the quantum signals mediating the feedback can preserve their coherence [17, 26]. Coherent feedback has been suggested both to support ground-state cooling in the resolved-sideband regime [27, 28, 29, 30] and to allow ground-state cooling in the unresolved-sideband regime [31], where we want to highlight a recent experimental realization [32].
@@ III.2 Feedback-assisted ground-state cooling
Figure 3: Plots of the final phonon number $n_{\mathrm{fin}}$ (on logarithmic scale) for the double-sided feedback scheme. (a) Final phonon number versus detuning $\Delta_{a}$ and phase accumulation $\phi$ with $\eta=0.95$ and $\kappa_{2}=\kappa_{1}$. (b) Final phonon number versus feedback efficiency $\eta$ with $\Delta_{a}/\Omega_{m}=1$, $\kappa_{2}=\kappa_{1}$, and two selected values of $\phi$ [with the orange solid and green dotted lines corresponding to the orange square and green circle in (a), respectively.] (c) Final phonon number versus decay ratio $\kappa_{2}/\kappa_{1}$ and feedback efficiency $\eta$ with $\Delta_{a}/\Omega_{m}=1$ and $\phi=0$. (d) Final phonon number versus decay ratio $\kappa_{2}/\kappa_{1}$ with $\Delta_{a}/\Omega_{m}=1$, $\phi=0$, and two selected values of $\eta$ [with the black dashed and red dotted lines corresponding to the same types of line in (c), respectively.] The white dashed line in (a) indicates $\Delta_{\text{eff}}=\Omega_{m}$ while the cyan areas in (b) and (d) represent the regimes of $n_{\mathrm{fin}}<1$. Other parameters are $\kappa_{1}/2\pi=0.25\,\text{MHz}$, $\Omega_{m}/2\pi=0.13\,\text{MHz}$, $\gamma_{m}/2\pi=0.12\,\text{Hz}$, $g_{0}/2\pi=50\,\text{Hz}$, $|\varepsilon_{\mathrm{p,eff}}|/2\pi=80\,\text{MHz}$, and $\bar{n}_{m}=9.6\times 10^{4}$ [53].
@@
We would like to point out that the effect of the double-sided coherent feedback is strongly dependent on the geometric symmetry of the cavity, i.e., the ratio $\kappa_{2}/\kappa_{1}$ between the decay rates at the two mirrors. As shown in Fig. 3(c), the minimum of $n_{\mathrm{fin}}$ always appears when $\kappa_{2}/\kappa_{1}\approx 1$. This can be understood again from the effective decay rate $\kappa_{\text{eff}}$ of the cavity, which is significantly reduced when $\kappa_{1}\approx\kappa_{2}$, $\text{mod}(\phi,2\pi)=0$, and $\eta\rightarrow 1$. Moreover, we focus in Fig. 3(d) on two selected cases ($\eta=0.8$ and $0.9$) where the system is always stable when $\kappa_{2}/\kappa_{1}$ varies. One can find that $n_{\mathrm{fin}}$ increases rapidly as the system deviates from the condition $\kappa_{2}/\kappa_{1}\sim 1$, hindering cooling. In view of this, the double-sided coherent feedback scheme is not a suitable candidate to facilitate ground-state cooling in the aforementioned Fano-mirror optomechanical setup, which by construction shows a highly asymmetric two-sided geometry. This conclusion is further verified in Appendix D, where we show that for $\kappa_{2}\ll\kappa_{1}$ such a double-sided coherent feedback can hardly affect the two optical normal modes.
@@ IV.2 Feedback-assisted ground-state cooling
Since the coherent feedback is introduced to the left Fano mirror in this case, we assume that the pumping field is applied to the right normal mirror. This allows the pumping field and the coherent feedback to be tuned independently, thus offering more flexibility for optimizing our cooling scheme. Specifically, this means that, in contrast to the previously shown two-sided feedback scheme, here the pumping amplitude $\varepsilon_{\mathrm{p}}$ is not modified by the feedback loop. The semiclassical steady-state values of the cavity and Fano modes are then
@@
Figure 7: (a) Linewidth $\kappa_{-}$ of the “-” normal mode as a function of reflection coefficient $r_{\text{CBS}}$ with $\phi/\pi=0$ and $\phi/\pi=1$. (b, c) Final phonon number $n_{\mathrm{fin}}$ (on logarithmic scale) versus reflection coefficient $r_{\mathrm{CBS}}$ and (b) phase accumulation $\phi$ and (c) extra efficiency factor $\eta_{\text{ex}}$. We assume $\eta_{\text{ex}}=0.9$ in (a) and (b) and $\phi=\pi$ in (c). Other parameters are identical to those in Fig. 5(a).
As mentioned above, the cooling effect can be greatly enhanced by the single-sided feedback scheme when $\phi=\pi$. This is quite different from the double-sided feedback scheme, where the system enters the effective sideband-resolved regime when $\phi=0$. The underlying reason can be understood from the interplay between the Fano resonance and the coherent feedback. The former primarily relies on strong enough optical dissipative coupling $\kappa_{1f}$, which can lead to a large linewidth splitting between the two normal modes (as also discussed in Appendix G). Importantly, the optical dissipative coupling would be significantly suppressed when $\phi=0$ due to the destructive interference arising from the coherent feedback, even if the feedback loop is not perfect ($\eta_{\text{ex}}<1$). This would cause the smaller normal-mode linewidth to increase, as shown in Fig. 7(a). In contrast, an appropriate constructive interference with $\phi=\pi$ can further reduce the smaller normal-mode linewidth.
@@ B.2 Single-sided feedback
$M^{\prime}V^{\prime}+V^{\prime}M^{\prime\text{T}}-N^{\prime}=0,$ (39)
@@ Appendix C Equipartition of energy in the ground-state cooling regimes
We first present an example for the double-sided feedback scheme in Figs. C.1(a), which correspond to the yellow solid line in Fig. 3(b). We find that the two mechanical variances are in good agreement, with their difference being negligible compared to their average values. For the single-sided feedback scheme, we show in Fig. C.1(c) that the difference between the two mechanical variances increases gradually with $\varepsilon_{\text{p}}$. This indicates that increasing the pumping amplitude is not always helpful for enhancing the cooling effect, even if it results in a lower final phonon number, as is shown in Fig. 5(c). In view of this, we choose an appropriate value for $\varepsilon_{\text{p}}$ in all other panels of Figs. 57, as indicated by the black dotted line in Fig. C.1(b), to ensure the equipartition of energy.
@@ Appendix D Optical normal modes with double-sided coherent feedback
Figure D.1: Effective resonance frequency $\omega_{-}$ and linewidth $\kappa_{-}$ of the “-” optical normal mode versus the feedback efficiency $\eta$ and phase accumulation $\phi$ for $\kappa_{2}/2\pi=0.6\,\text{GHz}$ [(a) and (b)] and $\kappa_{2}/2\pi=20\,\text{THz}$. All the panels share the same legend. Other parameters, except for those indicated in the panels, are $\kappa_{1}/2\pi=20\,\text{THz}$, $\kappa_{f}/2\pi=1.08\,\text{GHz}$, $\Delta_{a}/\Omega_{m}=30$, $\delta_{\Delta}/\Omega_{m}=18.2$, $\lambda/2\pi=805\,\text{MHz}$.
@@
As an example, Figs. D.1(a) and D.1(b) [Figs. D.1(c) and D.1(d)] show that the effective resonance frequency $\omega_{-}$ and linewidth $\kappa_{-}$ of the “-” normal mode are barely affected (significantly affected) by the coherent feedback when $\kappa_{2}\ll\kappa_{1}$ (when $\kappa_{2}=\kappa_{1}$). One can thus conclude that in the highly asymmetric case of $\kappa_{2}\ll\kappa_{1}$, the double-sided coherent feedback is inefficient for the Fano-mirror optomechanical setup.
@@ E.1 Standing-wave version
$H_{\text{loop}}=\sum_{j=1}^{j_{\text{max}}}\omega_{j}b_{j}^{{\dagger}}b_{j},$ (49)
@@
$\begin{split}H_{\text{int}}&=\left[\sum_{j=1}^{j_{\text{max}}}\xi_{j}a^{{\dagger}}\sin{(k_{j}d)}b_{j}+\text{H.c.}\right]+\sum_{j=1}^{j_{\text{max}}}H_{\text{om},j},\end{split}$ (50)
@@ Appendix G Comparison between Fano-mirror and coupled-cavity cooling scheme
Figure G.1: (a) Final phonon number $n_{\text{fin}}$ (on logarithmic scale) and mechanical variances ($\langle\delta x^{2}\rangle$ and $\langle\delta p^{2}\rangle$) versus detuning $\delta_{\Delta}$ for our Fano-mirror setup, with parameters differing from those considered elsewhere and in the absence of the single-sided coherent feedback (i.e., $\eta=0$). (b) Effective linewidths $\kappa_{\pm}$ (on logarithmic scale) of the optical normal modes versus the dimensionless control parameter $\zeta$ for $\eta=0$. We take $\xi=1$ in (a) and $\delta_{\Delta}=17.5$ in (b). The two panels share the same legend. Here we assume $\lambda/2\pi=0.9\,\text{GHz}$ and $\varepsilon_{\text{p}}/2\pi=0.8\,\text{THz}$. Other parameters are the same as those in Fig. 5(a).

### v3

@@ I Introduction
While the standard sideband cooling scheme [1] provides a powerful tool to access the mechanical ground state in optomechanical systems with high sideband resolution [14, 15, 16], in practice setups often are in the unresolved-sideband regime, meaning that simple sideband cooling no longer works. Quantum feedback has emerged as a candidate for ground-state cooling and control of optomechanical systems that are in the unresolved-sideband regime [17]. Compared to active feedback [18, 19, 20, 21, 22, 23, 24, 25], that has to face excess noise in the out-of-loop optical field and decoherence due to quantum measurement [22], coherent feedback is promising since the quantum signals mediating the feedback can preserve their coherence [17, 26]. Coherent feedback has been suggested both to facilitate ground-state cooling in the resolved-sideband regime [27, 28, 29, 30] and to allow ground-state cooling in the unresolved-sideband regime [31], where we want to highlight a recent experimental realization [32]. Ground-state cooling achieved by coherent feedback is, however, sensitive to limitations in the feedback efficiency, in particular for cavities with high losses.
@@ III.2 Feedback-assisted ground-state cooling
Figure 3: Plots of the final phonon number $n_{\mathrm{fin}}$ (on logarithmic scale) for the double-sided feedback scheme. (a) Final phonon number versus detuning $\Delta_{a}$ and phase accumulation $\phi$ with $\eta=0.95$ and $\kappa_{2}=\kappa_{1}$. (b) Final phonon number versus feedback efficiency $\eta$ with $\Delta_{a}/\Omega_{m}=1$, $\kappa_{2}=\kappa_{1}$, and two selected values of $\phi$ [with the orange solid and green dotted lines corresponding to the orange square and green circle in (a), respectively]. The orange curve approaches the minimum (i.e., $n_{\text{fin}}\approx 0.59$) for $\eta\approx 0.92$. (c) Final phonon number versus decay ratio $\kappa_{2}/\kappa_{1}$ and feedback efficiency $\eta$ with $\Delta_{a}/\Omega_{m}=1$ and $\phi=0$. (d) Final phonon number versus decay ratio $\kappa_{2}/\kappa_{1}$ with $\Delta_{a}/\Omega_{m}=1$, $\phi=0$, and two selected values of $\eta$ [with the black dashed and red dotted lines corresponding to the same types of line in (c), respectively.] The white dashed line in (a) indicates $\Delta_{\text{eff}}=\Omega_{m}$ while the cyan areas in (b) and (d) represent the regimes of $n_{\mathrm{fin}}<1$. Other parameters are $\kappa_{1}/2\pi=0.25\,\text{MHz}$, $\Omega_{m}/2\pi=0.13\,\text{MHz}$, $\gamma_{m}/2\pi=0.12\,\text{Hz}$, $g_{0}/2\pi=50\,\text{Hz}$, $|\varepsilon_{\mathrm{p,eff}}|/2\pi=80\,\text{MHz}$, and $\bar{n}_{m}=9.6\times 10^{4}$ [53].
@@
We would like to point out that the effect of the double-sided coherent feedback is strongly dependent on the geometric symmetry of the cavity, i.e., the ratio $\kappa_{2}/\kappa_{1}$ between the decay rates at the two mirrors. As shown in Fig. 3(c), the minimum of $n_{\mathrm{fin}}$ always appears when $\kappa_{2}/\kappa_{1}\approx 1$. This can be understood again from the effective decay rate $\kappa_{\text{tot,eff}}$ of the cavity, which is significantly reduced when $\kappa_{1}\approx\kappa_{2}$, $\text{mod}(\phi,2\pi)=0$, and $\eta\rightarrow 1$. Moreover, we focus in Fig. 3(d) on two selected cases ($\eta=0.8$ and $0.9$) where the system is always stable when $\kappa_{2}/\kappa_{1}$ varies. One can find that $n_{\mathrm{fin}}$ increases rapidly as the system deviates from the condition $\kappa_{2}/\kappa_{1}\sim 1$, showing how this asymmetry hinders ground-state cooling. Even with identical mirrors (i.e., $\kappa_{1}=\kappa_{2}$), the coherent feedback is further limited by its efficiency, especially for bad cavities. Specifically, to reach an effective sideband-resolved regime, the feedback efficiency must satisfy $1-\sqrt{\eta}<\Omega_{m}/\kappa_{\text{tot}}$, which is a stringent requirement for bad cavities (e.g., $\kappa_{\text{tot}}/\Omega_{m}\sim 10^{7}$ as will be considered in Sec. IV). In view of this, the double-sided coherent feedback scheme is not a suitable candidate to facilitate ground-state cooling in either the aforementioned Fano-mirror optomechanical setup, which by construction shows a highly asymmetric two-sided geometry, or conventional optomechanical setups with a poor sideband resolution $\Omega_{m}/\kappa_{\text{tot}}$. This conclusion is further verified in Appendix D, where we show that for $\kappa_{2}\ll\kappa_{1}$ such a double-sided coherent feedback can hardly affect the two optical normal modes.
@@ IV.2 Feedback-assisted ground-state cooling
Since the coherent feedback is introduced to the left Fano mirror in this case, we assume that the pumping field is applied to the right normal mirror. This allows the pumping field and the coherent feedback to be tuned independently, thus offering more flexibility for our cooling scheme. Specifically, this means that, in contrast to the previously shown two-sided feedback scheme, here the pumping amplitude $\varepsilon_{\mathrm{p}}$ is not modified by the feedback loop. The semiclassical steady-state values of the cavity and Fano modes are then
@@
Figure 7: (a) Linewidth $\kappa_{-}$ of the “–” normal mode as a function of reflection coefficient $r_{\text{CBS}}$ with $\phi/\pi=0$ and $\phi/\pi=1$. (b, c) Final phonon number $n_{\mathrm{fin}}$ (on logarithmic scale) versus reflection coefficient $r_{\mathrm{CBS}}$ and (b) phase accumulation $\phi$ and (c) extra efficiency factor $\eta_{\text{ex}}$. We assume $\eta_{\text{ex}}=0.9$ in (a) and (b) and $\phi=\pi$ in (c). Other parameters are identical to those in Fig. 5(a).
As mentioned above, the cooling effect in the Fano-mirror setup can be enhanced by the single-sided feedback scheme when $\phi=\pi$. This is quite different from the double-sided feedback scheme, where the standard optomechanical system (without the Fano mode) enters the effective sideband-resolved regime when $\phi=0$. This difference arises from the interplay between the Fano resonance and the coherent feedback. The former becomes highly effective with strong optical dissipative coupling $\kappa_{1f}$, which leads to a significant linewidth splitting between the two normal modes (as also discussed in Appendix G). Importantly, the optical dissipative coupling would be significantly suppressed when $\phi=0$, due to the destructive interference arising from the coherent feedback, even if the feedback loop is not perfect ($\eta_{\text{ex}}<1$). This would cause the smaller normal-mode linewidth to increase, as shown in Fig. 7(a). In contrast, by employing coherent feedback with constructive interference (e.g., $\phi=\pi$), the smaller normal-mode linewidth can be further reduced due to the effectively enhanced optical dissipative coupling.
@@
These results show that the combination of a Fano mirror with single-sided coherent feedback is a promising scheme to achieve ground-state cooling circumventing too strict requirements on the parameters of the Fano-mirror optomechanical setup. However, further optimization of parameters can lead to even lower phonon numbers, e.g., when increasing the pumping amplitude $\varepsilon_{\text{p}}$, reducing the optical coherent coupling $\lambda$, or increasing the cavity decay $\kappa_{1}$ through the Fano mirror.
@@ B.2 Single-sided feedback
$M^{\prime}V^{\prime}+V^{\prime}M^{\prime\text{T}}+N^{\prime}=0,$ (39)
@@ Appendix C Equipartition of energy in the ground-state cooling regimes
We first present an example for the double-sided feedback scheme in Fig. C.1(a), which corresponds to the orange solid line in Fig. 3(b). We find that the two mechanical variances are in good agreement, with their difference being negligible compared to their average values. For the single-sided feedback scheme, we show in Fig. C.1(c) that the difference between the two mechanical variances increases gradually with $\varepsilon_{\text{p}}$. This indicates that increasing the pumping amplitude is not always helpful for enhancing the cooling effect, even if it results in a lower final phonon number, as is shown in Fig. 5(c). In view of this, we choose an appropriate value for $\varepsilon_{\text{p}}$ in all other panels of Figs. 57, as indicated by the black dotted line in Fig. C.1(b), to ensure the equipartition of energy.
@@ Appendix D Optical normal modes with double-sided coherent feedback
Figure D.1: Effective resonance frequency $\omega_{-}$ and linewidth $\kappa_{-}$ of the “–” optical normal mode versus the feedback efficiency $\eta$ and phase accumulation $\phi$ for $\kappa_{2}/2\pi=0.6\,\text{GHz}$ [(a) and (b)] and $\kappa_{2}/2\pi=20\,\text{THz}$. All the panels share the same legend. Other parameters, except for those indicated in the panels, are $\kappa_{1}/2\pi=20\,\text{THz}$, $\kappa_{f}/2\pi=1.08\,\text{GHz}$, $\Delta_{a}/\Omega_{m}=30$, $\delta_{\Delta}/\Omega_{m}=18.2$, $\lambda/2\pi=805\,\text{MHz}$.
@@
As an example, Figs. D.1(a) and D.1(b) [Figs. D.1(c) and D.1(d)] show that the effective resonance frequency $\omega_{-}$ and linewidth $\kappa_{-}$ of the “–” normal mode are barely affected (significantly affected) by the coherent feedback when $\kappa_{2}\ll\kappa_{1}$ (when $\kappa_{2}=\kappa_{1}$). One can thus conclude that in the highly asymmetric case of $\kappa_{2}\ll\kappa_{1}$, the double-sided coherent feedback is inefficient for the Fano-mirror optomechanical setup.
@@ E.1 Standing-wave version
$H_{\text{loop}}=\sum_{j=1}^{j_{\text{max}}}\omega_{j}b_{j}^{{\dagger}}b_{j}$ (49)
@@
$\begin{split}H_{\text{int}}&=\left[\sum_{j=1}^{j_{\text{max}}}\xi_{j}a^{{\dagger}}\sin{(k_{j}d)}b_{j}+\text{H.c.}\right]+\sum_{j=1}^{j_{\text{max}}}H_{\text{om},j}\end{split}$ (50)
@@ Appendix G Comparison between Fano-mirror and coupled-cavity cooling scheme
Figure G.1: (a) Final phonon number $n_{\text{fin}}$ (on logarithmic scale) and mechanical variances ($\langle\delta x^{2}\rangle$ and $\langle\delta p^{2}\rangle$) versus detuning $\delta_{\Delta}$ for our Fano-mirror setup, with parameters differing from those considered elsewhere and in the absence of the single-sided coherent feedback (i.e., $\eta=0$). (b) Effective linewidths $\kappa_{\pm}$ (on logarithmic scale) of the optical normal modes versus the dimensionless control parameter $\zeta$ for $\eta=0$. We take $\zeta=1$ in (a) and $\delta_{\Delta}=17.5$ in (b). The two panels share the same legend. Here we assume $\lambda/2\pi=0.9\,\text{GHz}$ and $\varepsilon_{\text{p}}/2\pi=0.8\,\text{THz}$. Other parameters are the same as those in Fig. 5(a).